# Decoding Secret Memorization in Code LLMs Through Token-Level Characterization

### v1

@@ VI-E Case Study
Example 2 in Fig. 8(b) presents another scenario involving the configuration of multiple credentials, including a GOCI. HCR generates an implausible secret containing repeated “5”. While both DeSec and BS-5 generate plausible secrets, only the one generated by DeSec is real. This success of DeSec is due to the scoring model guided probability re-weighting enhancement, which increases the likelihood of selecting real secret tokens.
@@ VII Discussion
VII-A Threats to validity
Internally, the study relies on heuristic methods for feature selection, which may not capture all key factors influencing real secret generation. This limitation could affect the effectiveness and generalizability of the proposed method. Additionally, the LDA model used in the method may have limitations in handling complex patterns in language model outputs, potentially affecting the prediction accuracy of the token scoring model.
@@ VII-A Threats to validity
VII-B Implications
We summarize the implications of this work as follows:
  • Enhanced understanding of memorization in Code LLMs: Our study provides valuable insights into the memorization of sensitive information in Code LLMs, contributing to a better understanding of the internal workings of these models and the associated privacy risks.
  • Improved assessment of privacy leakage risks: By leveraging token-level features and guiding the decoding process, our method enables a more comprehensive assessment of the privacy leakage risks associated with Code LLMs, helping developers and users make informed decisions.
  • Advancement of responsible AI practices: Our work contributes to the advancement of responsible AI practices by highlighting the importance of addressing privacy concerns in Code LLMs and providing a framework for characterizing and extracting real secrets from them.
  • Stimulation of further research: Our study lays the foundation for further research in the area of memorization and privacy in Code LLMs, encouraging the development of more sophisticated techniques and theoretical frameworks to address these challenges.
@@ VIII Conclusion
In this paper, we present a novel approach to characterize and extract real secrets from Code LLMs based on token-level probabilities. Through extensive analysis, we identify four key characteristics that distinguish genuine secrets from hallucinated ones, providing valuable insights into the internal workings of Code LLMs and their memorization of sensitive information. To address the limitations of existing prompt engineering techniques, we propose DeSec, a two-stage method that leverages token-level features derived from the identified characteristics to guide the decoding process. Extensive experiments on five state-of-the-art Code LLMs and a diverse dataset demonstrate the superior performance of DeSec compared to existing baselines, achieving a significantly higher plausible rate and successfully extracting a larger number of real secrets from the victim models, enabling a more comprehensive assessment of the privacy leakage risks associated with Code LLMs.

### v2

@@ VI-F Case Study
Example 2 in Fig. 8(b) presents another scenario involving the configuration of multiple credentials, including a GOCI. HCR generates an implausible secret containing repeated “5”. While both DeSec and BS-5 generate plausible secrets, only the one generated by DeSec is real. This success of DeSec is due to the scoring model-guided probability re-weighting step enhancement, instead of letting LLM simply select tokens based on probability during the initial decoding process, which increases the likelihood of selecting real secret tokens.
@@ VII Discussion
This study sheds light on the memorization of sensitive information in Code LLMs, offering insights into their internal workings and privacy risks. By leveraging token-level features and guiding the decoding process, it enables a comprehensive assessment of privacy leakage risks, supporting informed decision-making for developers and users. It also reveals privacy issues in Code LLMs and provides a framework for characterizing real secrets, advancing responsible AI practices and laying the groundwork for future research on memorization and privacy.
VII-A Potential Mitigation
Secure Secret Management. Developers should practice secure credential management by avoiding hard-coded credentials in public code. Sensitive information should be stored in configuration files or managed with tools like HashiCorp Vault [56] and AWS Secrets Manager [57] that provide encrypted storage for secure access.
Training Data Decontamination. Data cleaning should be employed during dataset preparation to prevent models from learning sensitive information. In addition to regular expressions, tools like GitGuardian [58] and truffleHog [59] can further scan the data for sensitive content.
Privacy-Safe Model Training. Techniques such as differential privacy [60] can help mitigate privacy risks by introducing noise during training, which obscures individual data points and prevents sensitive information from being memorized and leaked by the model.
VII-B Threats to validity
Internally, the study relies on heuristic methods for feature selection, which may not capture all key factors influencing real secret generation. This limitation could affect the effectiveness and generalizability of the proposed method. Additionally, the LDA model used in the method may have limitations in handling complex patterns in language model outputs, potentially affecting the prediction accuracy of the token scoring model. Lastly, while we use both API validation and GitHub search, if a secret was once valid but has since expired and been deleted from GitHub, it may be misclassified as fake. So the number of real secrets represents a lower bound. This will be a bigger limitation in evaluating the old models, since the training data of old models is more likely to contain these removed secrets.
@@ VIII Conclusion
In this paper, we present a novel approach to characterize and extract real secrets from Code LLMs based on token-level probabilities. Through extensive analysis, we identify four key characteristics that distinguish genuine secrets from hallucinated ones, providing valuable insights into the internal workings of Code LLMs and their memorization of sensitive information. To address the limitations of existing prompt engineering techniques, we propose DeSec, a two-stage method that leverages token-level features derived from the identified characteristics to guide the decoding process. Extensive experiments on five state-of-the-art Code LLMs and a diverse dataset demonstrate the superior performance of DeSec compared to existing baselines, achieving a significantly higher plausible rate and successfully extracting a larger number of real secrets from the victim models, enabling a more comprehensive assessment of the privacy leakage risks associated with Code LLMs. The code and data have been made available in our replication package [46].
Acknowledgments
This work was supported by the National Key Research and Development Program of China (No. 2021YFB3101500), the National NSF of China (grants No.62302176, No.62072046, 62302181), the Key R&D Program of Hubei Province (2023BAB017, 2023BAB079), and the Knowledge Innovation Program of Wuhan-Basic Research (2022010801010083).